# Andreas Schlaginhaufen SYCAMORE, EPFL &Reda Ouhamma SYCAMORE, EPFL &Maryam Kamgarpour

### v1

@@ A.4 Miscellaneous
Denote $N_{T}=\sum_{t=1}^{T}\mathbbm{1}\left[{V_{t}\neq V_{t-1}}\right]$, and let $T^{\prime}<T$ be the last time the matrix $V_{t}$ was updated. Then,
$\det(V_{T})$ $\geq(1+C)/(1-1/e)\det(V_{T^{\prime}})$
$\geq((1+C)/(1-1/e))^{N_{T}}\det(\lambda I).$
The first inequality follows because $\det W_{T}\geq(1+C)\det V_{T^{\prime}}$ and because greedy D-Optimal design ensures that $\det V_{T}\geq\frac{1}{1-1/e}\det W_{T}$, where $W_{t}$ is defined in Algorithm 2.
Then, using the trace-determinant inequality, we have:
$\left({\frac{1+C}{1-1/e}}\right)^{N_{T}}\lambda^{d}\leq\left({\frac{\lambda d+T(LH_{\gamma})^{2}}{d}}\right)^{d}$
and then,
$N_{T}\leq\frac{d}{\log\left({\frac{1+C}{1-1/e}}\right)}\log\left({1+\frac{T(LH_{\gamma})^{2}}{\lambda d}}\right).$
We now present the proof for the regret bound, which proceeds similarly to that of Theorem 3.4 up to some modifications. The first change is in the regret decomposition, which needs to be adapted because we no longer compare to the past policy but rather to $\pi_{t}^{\prime}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\hat{\theta}_{t_{\operatorname{stop}}(t)},\varepsilon,\delta)$. The second change is using Lemma A.6 instead of Lemma A.5 to bound the sum of norms of trajectory features. Finally, the good events defined in the proof of Theorem 3.4 are slightly modified to account for the lazy design matrix and the new choice of comparator policy $\pi_{t}^{\prime}$.
@@
We work with the same two filtrations as before $\mathcal{F}_{t}=\sigma(x_{1},y_{1},\ldots,x_{t-1},y_{t-1})$ and $\mathcal{F}_{t}^{\theta}=\sigma(\mathcal{F}_{t},\tilde{\theta}_{t})$. And we recall that, given $\mathcal{F}_{t}$, $\tilde{\theta}_{t}$ in Algorithm 2 is sampled from $\mathcal{N}(\hat{\theta}_{t_{\operatorname{stop}}(t)},\beta_{t_{\operatorname{stop}}(t)}^{2}V_{t_{\operatorname{stop}}(t)}^{-1})$.
@@
    Let $C_{T}:=2\sqrt{\!T\!\left({\frac{8(1+C)d}{(1-1/e)^{2}}\log\left(\!1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\!\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{2}{\delta^{\prime}}}\right)$, and define $G_{2}$ as the event under which it holds that
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}\mid\mathcal{F}_{t}}\right]\leq C_{T},$
@@
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}\mid\mathcal{F}_{t}^{\theta}}\right]\leq C_{T}.$
@@
    Let $G_{3}=\left\{{\forall t\in[1,T]:V_{\tilde{\theta}_{t}}^{*}-V_{\tilde{\theta}_{t}}^{\pi_{t}}\leq\varepsilon}\right\}$, where $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$.
@@
We now compare each of these events to their counterparts in the proof of Theorem 3.4. The event $G_{1}$ is modified because we replace $V_{t}$ by $V_{t_{\operatorname{stop}}(t)}$ and $\hat{\theta}_{t}$ by $\hat{\theta}_{t_{\operatorname{stop}}(t)}$. The event $G_{1}$ still holds with probability at least $1-2\delta^{\prime}$ using the same concentration arguments as before. The event $G_{2}$ is also modified to account for the lazy design matrix, and it holds with probability $1-\delta^{\prime}$ thanks to Lemma A.6. Finally, the event $G_{3}$ remains unchanged.
@@
Step 2 (regret decomposition): Since Algorithm 2 uses a different comparator policy $\pi_{t}^{\prime}$ than Algorithm 1, we derive a new regret decomposition.
@@
where $C_{T}=2\sqrt{T\left({\frac{8(1+C)d}{(1-1/e)^{2}}\log\left(1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{2}{\delta^{\prime}}}\right)$.
For the first term in the regret decomposition, similarly to Appendix B, we have that:
@@
Step 3 (bounding ${r_{t}^{\text{TS}}}$): The proof for $\sum_{t}r_{t}^{\text{TS}}$ proceeds exactly like Appendix B up to Equation (5), this is because the probability of the optimism event $O_{t}$ and the events $E_{t}$ and $\hat{E}_{t}$ is unaffected by the change to the algorithm. Then, we have that:
$R^{\text{TS}}(T)=\sum_{t=1}^{T}r_{t}^{\text{TS}}\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]+T\varepsilon.$
@@
Step 4 (bounding ${r_{t}^{\text{MLE}}}$): This proof is similar to that of the first term in the regret decomposition. The only difference is that we use the confidence set for $\tilde{\theta}_{t}$ instead of $\hat{\theta}_{t_{\operatorname{stop}}(t)}$. Then, it holds that:
@@
$\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\sum_{t=1}^{T}\mathbb{E}[\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}^{\theta}],$

### v2

@@ Acknowledgments
Denote $N_{T}=\sum_{t=1}^{T}\mathbbm{1}\left[{V_{t+1}\neq V_{t}}\right]$, and let $t_{\operatorname{stop}}\leq T$ be the last time the matrix $V_{t}$ was updated. Then,
$\det(V_{T})$ $=\det(V_{t_{\operatorname{stop}}})\geq(1+C)^{N_{T}}\det(\lambda I).$
Here, we used that for two consecutive update rounds $t_{\operatorname{stop}}\geq t_{\operatorname{stop}}^{\prime}$ we have that $\det V_{t_{\operatorname{stop}}}\geq(1+C)\det V_{t_{\operatorname{stop}}^{\prime}}$. Then, using the trace-determinant inequality $\det A\leq\left({\operatorname{tr}(A)/d}\right)^{d}$, we have
$\left({1+C}\right)^{N_{T}}\lambda^{d}\leq\left({\frac{d\lambda+T(2LH_{\gamma})^{2}}{d}}\right)^{d},$
and
$N_{T}\leq\frac{d}{\log\left({1+C}\right)}\log\left({1+\frac{T(2LH_{\gamma})^{2}}{d\lambda}}\right).$
We now present the proof for the regret bound, which proceeds similarly to that of Theorem 3.2.1 up to some modifications. The first change is in the regret decomposition, which needs to be adapted because we no longer compare to the past policy but rather to $\pi_{t}^{\prime}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\hat{\theta}_{t_{\operatorname{stop}}(t)},\varepsilon,\delta)$. The second change is using Lemma D.5 instead of Lemma D.4 to bound the sum of norms of trajectory features. Finally, the good events defined in the proof of Theorem 3.2.1 are slightly modified to account for the lazy design matrix and the new choice of comparator policy $\pi_{t}^{\prime}$.
@@
We work with the same two filtrations as before $\mathcal{F}_{t-1}=\sigma(x_{1},y_{1},\ldots,x_{t-1},y_{t-1})$ and $\mathcal{F}_{t-1}^{\theta}=\sigma(\mathcal{F}_{t-1},\tilde{\theta}_{t})$. And we recall that, given $\mathcal{F}_{t-1}$, $\tilde{\theta}_{t}$ in Algorithm 3 is sampled from $\mathcal{N}(\hat{\theta}_{t_{\operatorname{stop}}(t)},\beta_{t_{\operatorname{stop}}(t)}^{2}V_{t_{\operatorname{stop}}(t)}^{-1})$.
@@
    Let $C_{T}:=2\sqrt{\dfrac{8(1+C)\,dT\,L^{2}H_{\gamma}^{2}}{\lambda}\log\!\Bigl(1+\frac{4TL^{2}H_{\gamma}^{2}}{\lambda d}\Bigr)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log(2/\delta)$, and define $G_{2}$ as the event under which it holds that
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}\mid\mathcal{F}_{t-1}}\right]\leq C_{T},$
@@
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t_{\operatorname{stop}}(t)}^{-1}}\mid\mathcal{F}_{t-1}^{\theta}}\right]\leq C_{T}.$
@@
    Let $A_{t}=\left\{{V_{\tilde{\theta}_{t}}^{*}-V_{\tilde{\theta}_{t}}^{\pi_{t}}\leq\varepsilon}\right\}$, where $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$, and let $G_{3}=\bigcap_{t=1}^{T}A_{t}$.
@@
We now compare each of these events to their counterparts in the proof of Theorem 3.2.1. The event $G_{1}$ is modified because we replace $V_{t}$ by $V_{t_{\operatorname{stop}}(t)}$ and $\hat{\theta}_{t}$ by $\hat{\theta}_{t_{\operatorname{stop}}(t)}$. The event $G_{1}$ still holds with probability at least $1-2\delta^{\prime}$ using the same concentration arguments as before. The event $G_{2}$ is also modified to account for the lazy design matrix, and it holds with probability $1-\delta^{\prime}$ thanks to Lemma D.5. Finally, the event $G_{3}$ remains unchanged.
@@
Step 2 (regret decomposition): Since Algorithm 3 uses a different comparator policy $\pi_{t}^{\prime}$ than Algorithm 1, we derive a new regret decomposition. On the good event $G$, we have
@@
where $C_{T}=2\sqrt{\dfrac{8(1+C)\,dT\,L^{2}H_{\gamma}^{2}}{\lambda}\log\!\Bigl(1+\frac{4TL^{2}H_{\gamma}^{2}}{\lambda d}\Bigr)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log(2/\delta)$.
For the first term in the regret decomposition, similarly to Appendix A, we have that:
@@
Step 3 (bounding $r_{t}^{\text{TS}}$}): The proof for $\sum_{t}r_{t}^{\text{TS}}$ proceeds exactly like Appendix A up to Equation (3), this is because the probability of the optimism event $O_{t}$ and the events $E_{t}$ and $\hat{E}_{t}$ is unaffected by the change to the algorithm. Then, we have that:
$R^{\text{TS}}(T)=\sum_{t=1}^{T}r_{t}^{\text{TS}}\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]+T\varepsilon.$
@@
Step 4 (bounding $r_{t}^{\text{MLE}}$}): This step is analogous to Appendix A. We have:
@@
$\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\sum_{t=1}^{T}\mathbb{E}[\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t-1}^{\theta}],$
@@
Appendix D Technical results
D.1 Confidence sequence
The first result is a confidence sequence for the maximum likelihood estimation. It is an elliptical relaxation of the likelihood ratio confidence sequence provided in Theorem 3.1 of lee2024unified.
[…]
Proof.
where $\mathcal{L}_{\mathcal{D}_{t}}$ was defined in Equation (2) and
$G_{t}(\hat{\theta}_{t},\theta^{*})$ $=\int_{0}^{1}(1-\tau)\left({\sum_{k=1}^{t-1}\dot{\sigma}(\langle\hat{\theta}_{t}+\tau(\theta^{*}-\hat{\theta}_{t}),x_{k}\rangle)x_{k}x_{k}^{\top}}\right)\mathop{}\!\mathrm{d}\tau$
$=\sum_{k=1}^{t-1}\left[{\int_{0}^{1}(1-\tau)\dot{\sigma}(\langle\hat{\theta}_{t}+\tau(\theta^{*}-\hat{\theta}_{t}),x_{k}\rangle)\mathop{}\!\mathrm{d}\tau}\right]x_{k}x_{k}^{\top}$
$\succeq\kappa^{-1}\sum_{k=1}^{t-1}x_{k}x_{k}^{\top}.$ (5)
Rearranging terms gives
$\stackrel{{\scriptstyle(ii)}}{{\geq}}(\theta^{*}-\hat{\theta}_{t})^{\top}\left({\kappa^{-1}\sum_{k=1}^{t-1}x_{k}x_{k}^{\top}}\right)(\theta^{*}-\hat{\theta}_{t})$
where $(i)$ follows from the first order optimality condition for $\hat{\theta}_{t}$ and $(ii)$ from the lower bound in (5). Rearranging again and applying Theorem 3.1 by lee2024unified for likelihood ratio confidence sequences with the Lipschitz constant of $\mathcal{L}_{\mathcal{D}_{t}}$ equal to $L_{t}=2LH_{\gamma}(t-1)$, we get
D.2 Optimism with constant probability
Lemma D.1 (Appendix A of [pmlr-v54-abeille17a]).
  1. 1.
  2. 2.
The anti-concentration property yields the following key result, which is required to prove a constant probability of optimism and subsequently control pessimism terms in the regret. While it has been proven by pmlr-v54-abeille17a in their linear Thompson sampling analysis, we provide a concise proof based on convex analysis for completeness.
Lemma D.2.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a continuous and convex function, and consider the ellipsoid $\mathcal{E}:=\left\{{\theta\in\mathbb{R}^{d}:\left\lVert\theta-\theta_{0}\right\rVert_{V}\leq b}\right\}$ for a positive definite matrix $V$ and $b>0$. If $\tilde{\theta}\sim\mathcal{N}(\theta_{0},b^{2}V^{-1})$, then $\Pr\left[{f(\tilde{\theta})\geq\max_{\theta\in\mathcal{E}}f(\theta)}\right]\geq 1/\left({4\sqrt{e\pi}}\right)$.
Proof.
Note that by definition of $\tilde{\theta}$ we have $\tilde{\theta}\stackrel{{\scriptstyle d}}{{=}}\theta_{0}+bV^{-1/2}\tilde{z}$, where $\tilde{z}\sim\mathcal{N}(0,I)$. Hence, considering $g(z):=f(\theta_{0}+bV^{-1/2}z)$, we have
$p:=\Pr\left[{f(\tilde{\theta})\geq\max_{\theta\in\mathcal{E}}f(\theta)}\right]=\Pr\left[{g(\tilde{z})\geq\max_{z\in\mathcal{B}^{d}(1)}g(z)}\right],$
where we used that $\theta_{0}+bV^{-1/2}z\in\mathcal{E}$ if and only if $z\in\mathcal{B}^{d}(1)$. Since $g$ is a continuous convex function, we can choose $\bar{z}\in\operatorname*{arg\,max}_{z\in\mathcal{B}^{d}(1)}g(z)$ such that $\left\lVert\bar{z}\right\rVert=1$. By optimality of $\bar{z}$ it holds that
$\partial g(\bar{z})\subseteq N_{\mathcal{B}^{d}(1)}(\bar{z}),$ (6)
where $N_{\mathcal{B}^{d}(1)}(\bar{z})=\left\{{h\in\mathbb{R}^{d}:0\geq\langle h,z-\bar{z}\rangle,\forall z\in\mathcal{B}^{d}(1)}\right\}=\left\{{h\in\mathbb{R}^{d}:h=\lambda\bar{z},\lambda\geq 0}\right\}$ denotes the normal cone to $\mathcal{B}^{d}(1)$ at $\bar{z}$. While this optimality condition for convex function maximization is somewhat standard (see e.g. [rockafellar1997convex, Theorem 32.4]), we can verify directly that by optimality of $\bar{z}$ we have
$\partial g(\bar{z})$ $:=\left\{{h\in\mathbb{R}^{d}:g(z)-g(\bar{z})\geq\langle h,z-\bar{z}\rangle,\forall z}\right\}$
$\subseteq\left\{{h\in\mathbb{R}^{d}:g(z)-g(\bar{z})\geq\langle h,z-\bar{z}\rangle,\forall z\in\mathcal{B}^{d}(1)}\right\}$
$\subseteq\left\{{h\in\mathbb{R}^{d}:0\geq\langle h,z-\bar{z}\rangle,\forall z\in\mathcal{B}^{d}(1)}\right\}=:N_{\mathcal{B}^{d}(1)}(\bar{z}).$
Since $g$ is convex and finite on all of $\mathbb{R}^{d}$, we have $\partial g(\bar{z})\neq\emptyset$. Therefore, the inclusion (6) implies that
[…]
In light of Lemma D.1, this establishes $p\geq 1/(4\sqrt{e\pi})$. ∎
The above Lemma is helpful in the following way: Let $\theta^{*}\in\mathcal{E}$, $\tilde{\theta}\sim\mathcal{N}(\theta_{0},b^{2}V^{-1})$, and $x_{\tilde{\theta}}=\operatorname*{arg\,max}_{x\in\mathcal{X}}\langle\tilde{\theta},x\rangle$ for some bounded subset $\mathcal{X}\subset\mathbb{R}^{d}$ (assuming the maximum exists). If $f$ is the support function of $\mathcal{X}$, i.e. $f(\theta)=\max_{x\in\mathcal{X}}\langle\theta,x\rangle$, then we have with probability at least $1/\left({4\sqrt{e\pi}}\right)$ that
$\max_{x\in\mathcal{X}}\langle\theta^{*},x\rangle=f(\theta^{*})\leq\max_{\theta\in\mathcal{E}}f(\theta)\leq f(\tilde{\theta})=\langle\tilde{\theta},x_{\tilde{\theta}}\rangle$.
Moreover, by part 2. of Lemma D.1, we still have that $\left\lVert\tilde{\theta}-\theta_{0}\right\rVert_{V}\leq\tilde{\mathcal{O}}\left({\sqrt{d}b}\right)$ with high probability. This idea will be key to the proofs of Theorems 3.2.1, 3.2.2, and 4.2.
D.3 Elliptical potential bounds
Another key component of the convergence proofs are the following so-called elliptical potential lemmas that provide an upper bound on the sum of norms of sequentially observed vectors in the norm induced by their design matrix.
Lemma D.3 (Lemma 19.4 of lattimore2020bandit).
Let $\left({x_{t}}\right)_{t\geq 1}\subset\mathbb{R}^{d}$ and for all $t\geq 1$, $\left\lVert x_{t}\right\rVert\leq L$, let $V_{t}=\lambda I+\sum_{k=1}^{t-1}x_{k}x_{k}^{\top}$ for some $\lambda>0$. Then,
In the analysis of our algorithms, a central challenge is to control the norms of policy feature differences $\phi(\pi_{t})-\phi(\pi_{t}^{\prime})$. However, the learner only observes the trajectory-level differences $x_{t}=\phi(\tau_{t})-\phi(\tau_{t}^{\prime})$, which are random variables with mean $\phi(\pi_{t})-\phi(\pi_{t}^{\prime})$. To overcome this, we build on Lemma D.3 and introduce new tools to bound the sum of norms of policy feature differences.
Lemma D.4 (Elliptical lemma).
Let $\{x_{t}\}_{t\geq 1}\subset\mathbb{R}^{d}$ be a sequence of random vectors adapted to a filtration $\{\mathcal{F}_{t}\}_{t\geq 1}$, and let $\|x_{t}\|\leq L$ almost surely for all $t\geq 1$. Let $V_{t}=\lambda I+\sum_{k=1}^{t-1}x_{k}x_{k}^{\top}$ for some $\lambda>0$. Then, the following holds:
1) Deterministic bound. For all $T\in\mathbb{N}$, almost surely,
$\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}^{2}\;\leq\;2d\log\!\Bigl(1+\frac{TL^{2}}{d\lambda}\Bigr)\;+\;\frac{3dL^{2}}{\lambda\log 2}\,\log\!\Bigl(1+\frac{L^{2}}{\lambda\log 2}\Bigr).$
2) High-probability bound. For all $\delta\in(0,1)$, with probability at least $1-\delta$,
$\forall T\in\mathbb{N},\quad\sum_{t=1}^{T}\mathbb{E}\!\big[\|x_{t}\|_{V_{t}^{-1}}\,\big|\,\mathcal{F}_{t-1}\big]$ $\leq 2\sqrt{T\left({2d\log\left(1\!+\!\frac{TL^{2}}{d\lambda}\right)\!\!+\!\frac{3dL^{2}}{\lambda\log 2}\log\left(1\!+\!\frac{L^{2}}{\lambda\log 2}\right)}\right)}$
$\quad+\frac{8L}{\sqrt{\lambda}}\log(1/\delta).$
The first statement is a small improvement over Lemma D.3 because it involves $\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}$ instead of $\min\{1,\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\}$ and maintains a similar upper bound. In the second statement, $\{x_{t}\}_{t\geq 1}$ represent trajectory features and their expected values are policy features. Hence, the second statement of the lemma above allows us to control the elliptical potentials of the policy features while only observing trajectory features.
Proof.
First statement: The proof of this result is based on the observation in [lattimore2020bandit, Exercise 19.3]. Namely, the number of times the term $\left\lVert x_{t}\right\rVert_{{V}_{t}^{-1}}^{2}$ can be larger than one is at most $\frac{3d}{\log(2)}\log(1+\frac{L^{2}}{\lambda\log(2)})$.
where the first term of the last inequality follows by definition of $\mathcal{T}_{T}$. The second term follows because the number of times $1\leq t\leq T$ not in $\mathcal{T}_{T}$ is at most $\frac{3d}{\log(2)}\log(1+\frac{L^{2}}{\lambda\log(2)})$ as previously discussed, and because $\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\leq L^{2}/\lambda$. Then, the proof is concluded by bounding the first sum on the right-hand side using Lemma D.3.
Second statement: The proof proceeds by using Lemma D.9. We have for any $\delta\in(0,1)$ that with probability $1-\delta$:
$\sum_{t=1}^{T}\mathbb{E}[\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t-1}]$ $\leq 2\sum_{t=1}^{T}\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}+\frac{8L}{\sqrt{\lambda}}\log(1/\delta)$
where the first inequality uses Lemma D.9, the second uses the Cauchy-Schwarz inequality, and the last follows from the first result of the lemma. ∎
We now present a variant of the elliptical potential lemma above, adapted for the case where the design matrix is updated with lazy updates and optimal design; see Algorithm 3 for more details.
Lemma D.5 (Lazy elliptical lemma).
Let $(x_{t})_{t\geq 1}\subset\mathbb{R}^{d}$ be a sequence of random vectors adapted to a filtration $(\mathcal{F}_{t})_{t\geq 1}$, with $\|x_{t}\|\leq L$ almost surely. Fix $\lambda\geq L^{2}$ and $C>0$.
Define $(V_{t},W_{t})_{t\geq 1}$ as in Algorithm 3:
  • $V_{1}=W_{1}=\lambda I$, $\mathcal{D}_{1}=\emptyset$, $t_{\operatorname{stop}}=1$.
  • For $t=1,\dots,T$:
    • If $\det(W_{t})>(1+C)\det(V_{t_{\operatorname{stop}}})$:
        Let $V_{t}=V_{t-1}+\sum_{x\in\mathcal{D}_{\text{opt}}}xx^{\top}$ be such that $\det V_{t}\geq\det W_{t}$ and $\left|{\mathcal{D}_{\text{opt}}}\right|\leq\left|{\mathcal{D}}\right|$.
        Set $t_{\operatorname{stop}}=t$, $\mathcal{D}=\emptyset$
    • Observe $x_{t}$ and append $\mathcal{D}=\mathcal{D}\cup\left\{{x_{t}}\right\}$.
    • Update
      $W_{t+1}=\begin{cases}V_{t}+x_{t}x_{t}^{\top},&\text{if }t=t_{\operatorname{stop}},\\ W_{t}+x_{t}x_{t}^{\top},&\text{otherwise,}\end{cases}\quad V_{t+1}=V_{t}.$
Then, the following holds:
1) Deterministic bound. For all $T\in\mathbb{N}$,
$\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}^{2}$ $\leq\dfrac{2(1+C)\,d\,L^{2}}{\lambda}\log\!\Bigl(1+\frac{TL^{2}}{\lambda d}\Bigr).$
2) High-probability bound. For a fixed $T\in\mathbb{N}$ and any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sum_{t=1}^{T}\mathbb{E}\bigl[\|x_{t}\|_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}\bigr]$ $\leq 2\sqrt{\dfrac{2(1+C)\,dT\,L^{2}}{\lambda}\log\!\Bigl(1+\frac{TL^{2}}{\lambda d}\Bigr)}+\frac{8L}{\sqrt{\lambda}}\log(1/\delta).$
Proof.
First statement: First note that $V_{t}\succeq\lambda I$ for all $t$, hence $\|x_{t}\|_{V_{t}^{-1}}^{2}\leq L^{2}/\lambda$. For non-update steps, $t\neq t_{\operatorname{stop}}$, we have $V_{t}\preceq W_{t}$ and $\det(W_{t})\leq(1+C)\det(V_{t})$. Hence, by Lemma D.8
$\forall x\in\mathbb{R}^{d},\quad\left\lVert x\right\rVert_{V_{t}^{-1}}^{2}\leq\frac{\det(V_{t}^{-1})}{\det(W_{t}^{-1})}\left\lVert x\right\rVert_{W_{t}^{-1}}^{2}\leq(1+C)\left\lVert x\right\rVert_{W_{t}^{-1}}^{2}.$ (7)
Using the inequality $\min\{1,z\}\leq 2\log(1+z)$ for $z\geq 0$, we get
$\min\{1,\|x_{t}\|_{V_{t}^{-1}}^{2}\}$ $\leq(1+C)\,\min\{1,\|x_{t}\|_{W_{t}^{-1}}^{2}\}\leq 2(1+C)\log\left({1+\|x_{t}\|_{W_{t}^{-1}}^{2}}\right)$
$=2(1+C)\log\frac{\det\left({W_{t+1}}\right)}{\det\left({W_{t}}\right)},$
where the last equality follows from $W_{t+1}=W_{t}+x_{t}x_{t}^{\top}$ and the matrix determinant lemma (see [abbasi2011improved, Lemma 11]).
For update steps, $t=t_{\operatorname{stop}}$, we have by construction $\det(V_{t})\geq\det(W_{t})$ and $W_{t+1}=V_{t}+x_{t}x_{t}^{\top}$. Therefore, with the same reasoning as above,
$\min\{1,\|x_{t}\|_{V_{t}^{-1}}^{2}\}\leq 2\log\left({1+\|x_{t}\|_{V_{t}^{-1}}^{2}}\right)=2\log\frac{\det\left({W_{t+1}}\right)}{\det\left({V_{t}}\right)}\leq 2\log\frac{\det\left({W_{t+1}}\right)}{\det\left({W_{t}}\right)}.$
By summing over $t$ and telescoping, we therefore get
$\sum_{t=1}^{T}\min\{1,\|x_{t}\|_{V_{t}^{-1}}^{2}\}\leq 2(1+C)\log\frac{\det\left({W_{T+1}}\right)}{\det\left({W_{1}}\right)}\leq 2(1+C)d\log\left({1+\dfrac{TL^{2}}{\lambda d}}\right).$
Using that $\left\lVert x_{t}\right\rVert_{V_{t}^{-1}}^{2}\leq(L^{2}/\lambda)\min\{1,\|x_{t}\|_{V_{t}^{-1}}^{2}\}$ concludes the proof of the first result.
Second statement: Analogously to Lemma D.4, the second statement follows from the first statement and Lemma D.9. ∎
Remark D.6.
Note that we have not used the trick from [lattimore2020bandit, Exercise 19.3] in the proof above, as it doesn’t trivially extend to our lazy setting.
D.4 Miscellaneous
Proposition D.7.
Proof.
Furthermore, any convex function is continuous over the relative interior of its effective domain $\operatorname{\operatorname{dom}}f=\left\{{x:f(x)<\infty}\right\}$ (see e.g. [rockafellar1997convex, Theorem 10.1]). Since $|f(\theta)|\leq\left\lVert\theta\right\rVert LH_{\gamma}$, this implies that $f$ is continuous over $\mathbb{R}^{d}$. ∎
Lemma D.8 (Lemma 12 of [abbasi2011improved]).
$\sup_{x\neq 0}\frac{x^{T}Ax}{x^{T}Bx}\leq\frac{\det(A)}{\det(B)}.$
This next lemma is one of the versions of Freedman’s inequality [freedman1975tail].
Lemma D.9 (Lemma 2 of zhu2022efficient).
Let $\left(z_{t}\right)_{t\leq T}$ be a real-valued sequence of random variables adapted to a filtration $\mathcal{F}_{t}$. If $0\leq z_{t}\leq B$ almost surely, then with probability at least $1-\delta$,
$\sum_{t=1}^{T}z_{t}\leq\frac{3}{2}\sum_{t=1}^{T}\mathbb{E}\left[z_{t}\mid\mathcal{F}_{t-1}\right]+4B\log\left(2\delta^{-1}\right)$
and
$\sum_{t=1}^{T}\mathbb{E}\left[z_{t}\mid\mathcal{F}_{t-1}\right]\leq 2\sum_{t=1}^{T}z_{t}+8B\log\left(2\delta^{-1}\right).$